OpenStudy (anonymous):

If f(x) = 2x2 + 3x and g(x) = x – 2, what is (f + g)(2)? A. 16 B. 14 C. 12 D. 10 D. 8

OpenStudy (welshfella):

to find (f + g)(x) just add the 2 functions the plug in x = 2

OpenStudy (welshfella):

* then

OpenStudy (anonymous):

How do you add the two functions. Sorry I've been away from school and I forgot how to do a lot of things.

OpenStudy (nino_jim):

Now you could have done f(2) + g(2) separately, which is: [2(2)2 + 3(2)] + [2 - 2] = 14 - 0 = 14

OpenStudy (anonymous):

So is that the answer?

OpenStudy (nino_jim):

Now replace x by 2, and get the answer", so that's 2(2)2 + 4(2) - 2 = 14

OpenStudy (nino_jim):

@RaisaD do you get it now

OpenStudy (nino_jim):

@RaisaD do you get it now

OpenStudy (anonymous):

Yes!

OpenStudy (nino_jim):

ok cool